# OctreeOcc: Efficient and Multi-Granularity Occupancy Prediction Using Octree Queries

### 09 Dec 2023 v2

@@ 4.4 Main Results
3D Semantic Scene Completion. To better assess the effectiveness of our approach, we conduct comparative experiments for the Semantic Scene Completion (SSC) task. As demonstrated in Tab 3, we compare our results with those of other SSC methods on the SemanticKITTI validation set. The outcomes indicate that, owing to the construction and correction of the octree, our model exhibits a more accurate perception of space. Our method outperforms other methods in the IoU metric, which evaluates the quality of reconstructed geometry. Additionally, for specific semantic classes such as car and vegetation, we achieve superior results, attributed to the enhanced treatment of objects facilitated by multi-granularity modeling.
4.5 Ablation Study & Analysis
In this subsection, we perform ablation studies on the Occ3d-nus validation set to assess the effectiveness of our proposed modules. Additionally, we analyze the impact of octree depth and the number of queries at each octree level.
Table 4: Ablation experiments of Modules on Occ3d-nuScenes dataset with reducing the input image size to 0.3x. “Sem.Init.” demotes semantic-guided octree initialization. “Iter.Rec.” means iterative structure rectification. The experiments were conducted on the NVIDIA A40 GPU.
@@ 4.5 Ablation Study & Analysis
$\checkmark$ 34.17 266 ms 27200M
$\checkmark$ $\checkmark$ 34.91 218 ms 18300M
$\checkmark$ $\checkmark$ $\checkmark$ 36.63 214 ms 17900M
$\checkmark$ $\checkmark$ $\checkmark$ 35.88 227 ms 18500M
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 37.40 224 ms 18500M
Figure 3: Illustration of octree structure rectification. The left figure displays the initially predicted octree structure, while the right figure depicts the octree structure after Iterative Structure Rectification module. It is evident that the predicted octree structure becomes more consistent with the object’s shape following the rectification module.
Effectiveness of Octree Queries. To validate the superiority of octree queries, we maintain consistent TSA and ICA settings while removing the proposed octree structure initialization and rectification modules. This allows for a comparison with the baseline that employs dense queries of size 100$\times$100$\times$16 (the largest size achievable within memory limitations). As illustrated in Table 4, the experimental results reveal that we continue to outperform with a 0.8 mIoU advantage, even with a memory saving of approximately 9G. This demonstrates that octree prediction can adeptly assign queries with different granularities to distinct semantic regions. The constructed octree queries are accommodate to various object shapes, showcasing a more efficient utilization of computational budget.
Table 5: Ablation for different octree depth on Occ3d-nuScenes $val$ set with reducing the input image size to 0.3x. The experiments were conducted on the NVIDIA A40 GPU.
Query Form Query Resolution mIoU Latency Memory
(a) 3D Voxel 50$\times$50$\times$4 28.51 129 ms 7400M
@@
(d) 100$\times$100$\times$16 34.10 266 ms 27200M
(e) Octree Query 50$\times$50$\times$4 to 100$\times$100$\times$8 32.02 182 ms 12400M
(f) 50$\times$50$\times$4 to 100$\times$100$\times$16 33.76 207 ms 16800M
(g) 50$\times$50$\times$4 to 200$\times$200$\times$16 37.40 224 ms 18500M
Table 6: Ablation for the choice of query selection ratio on Occ3d-nuScenes $val$ set with reducing the input image size to 0.3x. The experiments were conducted on the NVIDIA A40 GPU.
Query Form Query Selection Ratio mIoU Latency Memory
(a)) Octree Query 25%, 50% 36.73 220 ms 18000M
(b) Octree Query 25%, 60% 36.12 255 ms 21000M
(c) Octree Query 20%, 60% 37.10 224 ms 18500M
Figure 4: Qualitative results on Occ3D-nuScenes validation set. The first row displays input multi-view images, while the second row showcases the occupancy prediction results of PanoOcc[41], FBOCC[19], our methods, and the ground truth.
Effectiveness of Semantic-guided Structure Initialization module. To underscore the significance of the initial octree structure, we exclude the Semantic-guided Octree Initialization module and employ a MLP to predict the octree structure from randomly initialized queries. This results in a performance decrease of 1.7 mIoU, highlighting that despite having ground truth for supervision, the structurally predicted information from the initialized query is inaccurate due to the absence of coding for obtaining valid information. The incorporation of semantic priors proves crucial as it serves as a valuable reference to enhance the quality of the initialized octree, thereby improving overall model performance.
Effectiveness of Iteritive Structure Rectification module. We additionally perform an ablation study on the Iterative Structure Rectification module, as shown in Table 4. The incorporation of this module has led to noticeable improvements in performance. Additionally, we have presented a visualization of the octree structure both before and after correction (Fig. 3). These visualizations highlight the module’s role in rectifying inaccuracies in the initialized octree structure, thereby enhancing its compatibility with input scenarios. Consequently, this refinement contributes to the efficiency of octree query expression, positively impacting overall performance.
Discussion on the depth of octree. As shown in Tab. 5, we conducted experiments to investigate the impact of variations in the depth of the constructed octree. (a), (b), (c), and (d) illustrate the performance of four different sizes of 3D queries, while (e), (f), and (g) depict the performance of octree queries with three different depths and initial query sizes. By comparing (e) with (c) and (f) with (d), it is evident that our octree query approach achieves performance comparable to dense queries while significantly reducing memory consumption. (g) demonstrates that we can further refine modeling granularity, surpassing the constraints imposed by the excessive graphics memory overhead (>80G) associated with the 200$\times$200$\times$16 dense modeling and acheving higher performance.
Discussion on query selection ratio of each level in the octree. In Table  6, we present the results obtained by maintaining various query ratios across different levels of the octree. A comparison between (a) and (b) reveals that an excessive number of fine-grained queries leads to a degradation in performance. This is primarily due to the fact that the scenario does not predominantly consist of foreground points requiring modeling at such a detailed level. Further comparison between (a) and (c) demonstrates that, given the prevalence of coarsest-grained semantic regions, particularly those containing mostly empty spaces, modeling this portion at a fine-grained level adversely impacts computational efficiency. Consequently, determining the optimal number of queries with different granularities should align with the statistical distribution of object granularity within the scene. This ensures that queries of varying granularity can effectively handle semantic regions corresponding to their specific granularity levels.
4.6 Visualization
The Fig. 4 displays qualitative results obtained by our methods and other occupancy prediction solutions on the Occ3D-nus validation set, where the resolution of the voxel predictions is 200$\times$200$\times$16. The figure illustrates that our approach comprehensively understands the structure of the scene, showcasing superior performance in scene understanding.
@@ 5 Conclusions
In conclusion, our paper introduces OctreeOcc, a novel 3D occupancy prediction framework that addresses the limitations of traditional dense-grid representations in understanding 3D scenes. OctreeOcc’s adaptive utilization of octree representations enables the capture of valuable information with variable granularity, catering to objects of diverse sizes and complexities. Our extensive experimental results affirm OctreeOcc’s capability to attain state-of-the-art performance in 3D occupancy prediction while concurrently reducing computational overhead.
Supplementary Material
Appendix A More Details
@@
For Semantic-Guided Octree Initialization, we started by obtaining semantic segmentation labels for images through the projection of occupancy labels onto the surround-view image. Subsequently, a UNet semantic segmentation model was trained using these labels. To initialize the octree structure and queries, we conducted offline generation of semantic segmentation result graphs. This offline approach helps avoid unfair comparisons that may arise from simultaneous training with the occupancy prediction model. The initialization process involves the random initialization of dense queries and the coordinates of each query’s center point serve as reference points, which are projected onto the range-view image. If projected onto a ground pixel (representing driveable surface, other flat, or sidewalk), the probability increases by 0.1. Conversely, if projected onto a background pixel (excluding ground classes), the probability increases by 0.5. If a reference point is projected onto a foreground pixel, the probability of requiring a split at that position increases by 1.0. This process assigns a split probability to each query, and the octree mask is then constructed through average pooling, capturing split probabilities at different query levels. After obtaining the octree mask, we designate the top 20% confidence queries as parent queries in level 1, while the remaining queries become Leaf queries and remain unsplit. Moving to level 2, after splitting the parent queries into octants, the top 60% confidence positions are selected as new parent queries, and the remainder as leaf queries. By storing leaf queries at each level, we construct a sparse and multi-granularity octree structure for queries.
@@
For Iterative Structure Rectification, in level 1, we retain predictions for the top 10% of positions with confidence. For the remaining positions, we use a 2-layer MLP to predict probabilities. The new probabilities are then calculated by blending them with the existing probabilities, with a weight distribution of 60% for the new probabilities and 40% for the old ones. We identify the top 10% of positions with the new probability values as the required splits, constituting the structure of the new level 1. Similarly, in level 2, we preserve predictions for the top 30% of positions with confidence. For the positions not in the top 30%, we predict probabilities using a 2-layer MLP. The new probabilities are computed by combining them with the original probabilities, with a weight distribution of 50% for each. The top 30% of the new probability values are then selected as the positions that need to be split, forming the structure of the new level 2.
Appendix B Comparison of efficiency on the SemanticKITTI dataset
As shown in Tab. 7, we evaluate the efficacy of our approach by comparing it to alternative methods on the SemanticKITTI dataset. The results demonstrate that our approach not only outperforms these methods in terms of performance but also requires significantly less memory and exhibits lower latency in comparison. It is evident that employing sparse octree queries effectively mitigates computational overhead while maintaining robust performance.
Table 7: Comparison of query form and efficiency on the SemanticKITTI $val$ set. The experiments were conducted on the NVIDIA A40 GPU.
Methods IoU mIoU Latency Memory
TPVFormer[9] 35,61 11,36 179 ms 23000M
VoxFormer[17] 44.02 12.35 177 ms 23200M
OccFormer[45] 36.50 13.46 173 ms 22400M
Symphonics[10] 41.44 13.44 187 ms 22000M
Ours 44.71 13.12 162 ms 19000M
Appendix C Discussion on Octree Structure Quality
In Tab. 8, we present an analysis of the influence of the proposed initialization and correction modules on the octree’s quality. It is observed that continuous correction of the octree structure throughout the network learning process resulted in the rectification of several inaccurate predictions (approximately a 10% increase in mIoU predicted at each level).
Table 8: Comparison of octree quality at different stages
Stage mIoU(L1) mIoU(L2)
Initialized 57.34 51.28
After the 1st Rectification 60.13 53.95
After the 2nd Rectification 62.27 56.79
@@ Appendix C Discussion on Octree Structure Quality
Appendix D More Visualization

### 19 Mar 2024 v3

@@ 4.4 Results
3D Semantic Scene Completion. To better evaluate the effectiveness of our approach, we conduct comparative experiments for the Semantic Scene Completion (SSC) task. As demonstrated in Tab 2, we compare our results with those of other SSC methods on the SemanticKITTI validation set. Our model demonstrates a more accurate perception of space due to octree construction and correction, outperforming others in IoU metric for geometry reconstruction. Additionally, for specific semantic classes such as car and vegetation, we achieve superior results, attributed to the enhanced treatment of objects facilitated by multi-granularity modeling. Additionally, Tab 3 also indicates that our method consumes fewer computational resources than other dense query-based methods.
Table 3: Comparison of query form and efficiency with SOTA methods on the Occ3D-nuScenes (left table) and SemanticKITTI (right table) datasset.
Methods Query Form mIoU Latency Memory
BEVFormer[21] 2D BEV 23.67 302ms 25100M
TPVFormer[11] 2D tri-plane 28.34 341ms 29000M
PanoOcc[50] 3D voxel 42.13 502ms 35000M
FBOCC[22] 3D voxels & 2D BEV 43.41 463ms 31000M
Ours Octree Query 44.02 386ms 26500M
Methods Query Form IoU mIoU Latency Memory
TPVFormer[11] 2D tri-plane 35.61 11.36 179ms 23000M
OccFormer[56] 3D voxel 36.50 13.46 172ms 22400M
VoxFormer[18] 3D voxel 44.02 12.35 177ms 23200M
Symphonics[12] 3D voxel 41.44 13.44 187ms 22000M
Ours Octree Query 44.71 13.12 162ms 19000M
Table 4: Ablation experiments of Modules on Occ3d-nuScenes dataset.
Table 5: Comparison of octree structure quality at different stages.
@@
(a) $\checkmark$ 34.17 266 ms 27200M
(b) $\checkmark$ $\checkmark$ 34.91 218 ms 18300M
(c) $\checkmark$ $\checkmark$ $\checkmark$ 36.63 214 ms 17900M
(d) $\checkmark$ $\checkmark$ $\checkmark$ 35.88 227 ms 18500M
(e) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 37.40 224 ms 18500M
Stage mIoU
level 1 to 2 level 2 to 3
(a) Initialized w/o unbalanced assignment 45.79 33.60
Initialized w. unbalanced assignment 57.34 51.28
(b) After the 1st Rectification 60.13 53.95
After the 2nd Rectification 62.27 56.79
Table 5: Comparison of octree structure quality at different stages.
4.5 Ablation Study
In this subsection, we perform ablation studies on the Occ3d-nus validation set to assess the effectiveness of our proposed modules. All the experiments are conducted on the NVIDIA A40 GPU with reducing the input image size to 0.3x.
Effectiveness of Octree Queries. To validate the superiority of octree queries, we maintained consistent TSA and ICA settings while removing the proposed octree structure initialization and rectification modules. This facilitated a comparison with the baseline employing dense queries of size 100$\times$100$\times$16 (the largest size achievable within memory limitations). As illustrated in Tab. 5 (b), experimental results consistently demonstrate our outperformance, with a notable 0.8 mIoU advantage, despite achieving a memory saving of approximately 9G. This underscores the adeptness of octree prediction in allocating queries with varying granularities to diverse semantic regions. Moreover, the constructed octree queries exhibit adaptability to various object shapes, thereby optimizing the utilization of computational resources.
Effectiveness of Semantic-guided Structure Initialization module. To underscore the significance of the initial octree structure, we exclude the Semantic-guided Octree Initialization module and employ a MLP to predict the octree structure from randomly initialized queries. This results in a 1.7 mIoU performance decrease, highlighting the inaccuracy of structurally predicted information from the initialized query due to the absence of valid information coding. Incorporating semantic priors proves crucial as they enhance the quality of the initialized octree, thereby improving overall model performance. Meanwhile, Tab. 5(a) evaluates the effectiveness of the initial octree structure. We can observe that assigning different initial weights to voxels projected onto different semantic regions directs the initial structure to focus more on the effective areas of the scene, thus improving the quality of the octree structure.
Effectiveness of Iteritive Structure Rectification module. We additionally perform an ablation study on the Iterative Structure Rectification module, as shown in Tab. 5(d). The incorporation of this module has led to noticeable improvements in performance. Meanwhile, Tab. 5(b) shows this rectification gradually rectifies areas where structural predictions are incorrect. Consequently, this refinement contributes to the efficiency of octree query expression, positively impacting overall performance.
4.6 More Analysis
In this subsection, we provide more analysis experiments on the Occ3d-nus validation set. All the experiments are conducted on the NVIDIA A40 GPU with reducing the input image size to 0.3x.
Discussion on the depth of octree. Tab. 6 presents experiments on octree depth variations. (a)-(d) show performance with varying 3D query sizes, while (e)-(h) depict octree query performance with different depths and initial resolutions. Comparing (e) to (c) and (f) to (d) reveals our approach achieves comparable performance to dense queries, significantly reducing resource consumption. (g) shows that the predetermined octree depth should not be excessive. While reducing memory usage, the imperfect predictions of octree splitting result in accumulated errors, leading to performance degradation as the depth increases.
Table 6: Ablation for different octree depth on Occ3d-nuScenes $val$ set.
Query Form Octree Depth Query Resolution mIoU Latency Memory
(a) 3D voxel N/A 50$\times$50$\times$4 28.51 129ms 7400M
@@ 4.6 More Analysis
(d) 100$\times$100$\times$16 34.17 266ms 27200M
(e) Octree Query 2 50$\times$50$\times$4/100$\times$100$\times$8 32.02 182ms 12400M
(f) 3 50$\times$50$\times$4/100$\times$100$\times$8/100$\times$100$\times$16 33.76 207ms 16800M
(g) 4 25$\times$25$\times$2/50$\times$50$\times$4/100$\times$100$\times$8/100$\times$100$\times$16 34.88 193 ms 15800M
(h) 3 50$\times$50$\times$4/100$\times$100$\times$8/200$\times$200$\times$16 37.40 224ms 18500M
Table 7: Ablation for the choice of query selection ratio on Occ3d-nuScenes $val$ set.
Table 8: Comparison with another octree method.
Selection Ratio mIoU Latency Memory
(a) 10%, 60% 34.47 191ms 14500M
(b) 15%, 60% 35.01 203ms 16300M
(c) 25%, 50% 36.73 220ms 18000M
(d) 25%, 60% 36.12 255ms 21000M
(e) 20%, 60% 37.40 224ms 18500M
mIoU Latency Memory
baseline 34.10 266 ms 27200M
OGN[42] 33.39 212 ms 24300M
Ours 37.40 224 ms 18500M
Table 8: Comparison with another octree method.
Discussion on query selection ratio of each level in the octree. In Tab. 8, we present results from maintaining different query ratios in various octree level. Results show that too few or too many queries degrade performance. Inadequate queries result in an imperfect scene representation, especially for detailed regions (a,b vs e). Conversely, excessive queries impact computational efficiency, particularly for coarse-grained regions with empty spaces (c vs d and c,d vs e). Optimizing query numbers based on scene object granularity distribution ensures effective handling of semantic regions at different levels.
Analysis of Various Usage of Octree. As a classic technique, octree is employed in various tasks [42, 7, 15, 41, 46]. Despite differences in addressed problems, we compare our method with OGN[42], which proposes an octree-based upsampling approach. We keep the similar setup in Tab. 5 (b) but substitute the deconvolution decoder with OGN’s octree decoder. Results in Tab. 8 indicate that employing octree solely in the decoder fails to mitigate excessive computational costs and yields sub-optimal performance, mainly due to the high query count during encoding.
@@ 5 Conclusions
In conclusion, our paper introduces OctreeOcc, a novel 3D occupancy prediction framework that addresses the limitations of dense-grid representations in understanding 3D scenes. OctreeOcc’s adaptive utilization of octree representations enables the capture of valuable information with variable granularity, catering to objects of diverse sizes and complexities. Our extensive experimental results affirm OctreeOcc’s capability to attain state-of-the-art performance in 3D occupancy prediction while concurrently reducing computational overhead.
Appendix
Appendix 0.A More Details
@@
For the Semantic-Guided Octree Initialization, our approach commences with acquiring semantic segmentation labels for images by projecting occupancy labels onto the surround-view images. Subsequently, a UNet is trained using these labels. The initialization process entails randomly initializing dense queries, where each query’s center point serves as a reference point projected onto the range-view images. If a reference point is projected onto a ground pixel ($i.e.$, driveable surface, other flat, or sidewalk), the probability increases by 0.1. Conversely, if projected onto a background pixel (excluding ground classes), the probability increases by 0.5. Projection onto a foreground pixel increases the probability of requiring a split at that position by 1.0. This process assigns a split probability to each query, and the octree mask is constructed through average pooling, capturing split probabilities at different query levels. After obtaining the octree mask, we designate the top 20% confidence queries as parent queries in level 1, while the remaining queries become Leaf queries and remain unsplit. Moving to level 2, after splitting the parent queries into octants, the top 60% confidence positions are selected as new parent queries, and the remainder as leaf queries. By storing leaf queries at each level, we construct a sparse and multi-granularity octree structure for queries.
@@
In Iterative Structure Rectification, at level 1, we retain predictions for the top 10% of positions with confidence. For the remaining positions, a 2-layer MLP is utilized to predict probabilities. These new probabilities are blended with the existing probabilities, with a weight distribution of 60% for the new probabilities and 40% for the old ones. The top 10% of positions with the new probability values are identified as the required splits, shaping the structure of the new level 1. Similarly, at level 2, predictions for the top 30% of positions with confidence are preserved. For positions not in the top 30%, probabilities are predicted using a 2-layer MLP. The new probabilities are computed by merging them with the original probabilities, with an even weight distribution of 50% for each. The top 30% of the new probability values are then selected as the positions necessitating splitting, delineating the structure of the new level 2.
Appendix 0.B Octree node index calculation
The hierarchical structure of the octree, particularly the assignment of queries to respective levels, is determined based on the octree mask $M_{o}$ and the query selection ratio denoted as $\alpha=\{\alpha^{1},\alpha^{2},\ldots,\alpha^{L-1}\}$. These ratios govern the number of subdivisions at each level, thereby defining the hierarchical organization of the octree. The procedure is as follows. For level l, queries with $M_{o}^{l}$ values within the top $\alpha^{l}$ percentile are identified as candidates for octant subdivision. Subsequently, within octants that have undergone one previous subdivision at the next level, queries are once again selected based on their values falling within the top $\alpha^{2}$ percentile, initiating another round of subdivision. This process continues iteratively until reaching the final level of the octree.
Simultaneously, exploiting the octree structure facilitates the direct conversion of sparse octree queries into dense queries to align with the desired output shape. For a query $q_{octree}$ at level $l$ with the index $(a,b,c)$, the indexes of its corresponding $8^{L-l}$ children nodes in level $L$ are determined by $(a\times 2^{L-l}+a_{offset},b\times 2^{L-l}+b_{offset},c\times 2^{L-l}+c_{offset})$, where $a_{offset}$, $b_{offset}$, and $c_{offset}$ are independent, ranging from 00 to $2^{L-l}$. Here, $L$ denotes the depth of the octree. During this process, we allocate the feature of $q_{octree}$ to all of these positions. By iteratively applying this procedure to all queries at each level, we effectively transform $Q_{octree}$ into $Q_{dense}$.
Appendix 0.C More discussion of octree initialization
Given that the FloSP method outlined in MonoScene[3] incorporates a 3D to 2D projection operation, similar to our initialization approach, we additionally adapted this method for comparison. Specifically, we employed FloSP to extract 3D voxel features from 2D image features. Subsequently, we applied a Multi-Layer Perceptron (MLP) to predict the splitting probability of each voxel, replacing the randomly initialized queries used in the original ablation experiments. The results indicate that, although this operation outperforms predictions from randomly initialized queries, it is still constrained by insufficient information, resulting in a decline in overall performance.
Table 9: More ablation of octree initialization
Initialization Method mIoU
(a)) Randomly initialised queries 34.91
(b) Voxel features from FLoSP 35.72
(c) Semantic-Guided Octree Initialization 37.40
Appendix 0.D More Visualization
Fig. 6 illustrates the results of occupancy prediction alongside the corresponding octree structure. For clarity in visualization, we employ distinct colors to represent voxels at various levels of the octree prediction, based on their occupancy status. For improved visualization, only a portion correctly corresponding to the occupancy prediction is displayed, rather than the entire octree structure, ensuring clarity and focus on the relevant information. Level 3 (voxel size: 0.4m $\times$ 0.4m $\times$ 0.4m) is depicted in light gray, level 2 (voxel size: 0.8m $\times$ 0.8m $\times$ 0.8m) in medium gray, and level 1 (voxel size: 1.6m $\times$ 1.6m $\times$ 1.6m) in dark gray. It’s worth noting that level 1 voxels, predominantly situated in free space and within objects, might be less intuitively discernible. Nonetheless, this image underscores the efficacy of octree modeling, which tailors voxel sizes to different semantic regions, enhancing representation accuracy.
@@ Appendix 0.D More Visualization
Figure 6: Visulization of octree structure. The first row displays input multi-view images, while the second and third rows showcase the occupancy prediction results and the corresponding octree structure prediction results.